cathrinerose:

If f(x) = 3 – 2x and mc019-1.jpg, what is the value of mc019-2.jpg?

cathrinerose:

what yall think

Shadow:

You'll need to post a picture of the file or upload the file, or write out what the file is saying. I don't know what "mc019-1.jpg" is lol

cathrinerose:

its f(x)=3-2x and g(x)=1/x+5 what is the value of (f/g)(8)

cathrinerose:

here u go

Shadow:

\[\frac{ 3 - 2(8) }{ \frac{ 1 }{ 8 + 5}}\]

Shadow:

If you can't see, that's 1/(8 + 5)

cathrinerose:

thats the answer

Shadow:

Well, you are dividing the expression of f by the expression of g, when x = 8

Shadow:

You need to simplify what I wrote, using PEMDAS, as I basically just wrote out that^